# Latent Diffusion Model Based Denoising Receiver for 6G Semantic Communication: From Stochastic Differential Theory to Application

### v1

@@ III System Model and Problem Formulation
The end-to-end system comprises a source, a joint source–channel encoder, an additive white Gaussian noise (AWGN) channel, a denoising module, and a decoder. Let $\bm{x}\in\mathbb{R}^{n}$ denote the original source data. The source encoder, parameterized by $\bm{\phi}$, maps $\bm{x}$ into a compact semantic latent vector $\bm{z}\in\mathbb{R}^{d}$, with $d\ll n$, via
@@
where $f_{\bm{\phi}}:\mathbb{R}^{n}\to\mathbb{R}^{d}$ is typically realized by a variational autoencoder (VAE). The semantic latent vector $\bm{z}$ is transmitted over an AWGN channel, yielding the received representation
$\tilde{\bm{z}}=\bm{z}+\bm{n},\quad\bm{n}\sim\mathcal{N}(0,\sigma^{2}\bm{I}),$ (7)
where $\sigma^{2}$ denotes the channel noise power.
At the receiver, the corrupted latent variable $\tilde{\bm{z}}$ undergoes denoising to mitigate the impact of channel noise. Let $g_{\bm{\psi}}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ represent the denoising function parameterized by $\bm{\psi}$, and $h_{\bm{\theta}}:\mathbb{R}^{d}\to\mathbb{R}^{n}$ denote the semantic decoder parameterized by $\bm{\theta}$. The final reconstructed data $\hat{\bm{x}}$ is obtained as
@@
$\min_{\bm{\phi},\bm{\psi},\bm{\theta}},\mathbb{E}_{\bm{x},\bm{n}}\left[\left\|\bm{x}-h_{\bm{\theta}}\left(g_{\bm{\psi}}(\bm{z}+\bm{n})\right)\right\|^{2}\right],$ (9)
subject to the encoder constraint $\bm{z}=f_{\bm{\phi}}(\bm{x})$. This formulation captures the full stochasticity of the channel and reflects the end-to-end performance of the semantic transmission process. Two primary challenges arise in solving this problem. First, due to the dimensionality reduction $d\ll n$, the decoder must reconstruct a high-dimensional signal from a compressed and noise-contaminated latent representation. This makes the inverse mapping severely ill-posed and sensitive to noise perturbations. Second, the Gaussian noise introduced during transmission alters the distribution of the latent variable at the receiver, causing a significant mismatch between the training and inference distributions, commonly referred to as the OOD problem. These challenges underscore the need for a robust, distribution-aware denoising mechanism that can generalize across varying SNRs and input distributions. To address this, we propose a theoretically grounded denoising approach based on stochastic differential aligns and generative diffusion models, which will be detailed in the next section.
@@ IV-A Theoretical Basis of Diffusion Model-Based Denoising
$d\bm{x}_{t}=f_{t}\bm{x}_{t},dt+g_{t},d\bm{w}_{t},$ (10)
where $\bm{x}_{t}\in\mathbb{R}^{d}$ denotes the latent variable at time $t\in[0,T]$, $\bm{w}_{t}$ is a standard Wiener process, $f_{t}$ is a time-dependent drift coefficient, and $g_{t}$ is the diffusion coefficient. For tractability, we set $g_{t}=1$ and define the drift as $f_{t}\bm{x}_{t}=\bm{h}_{t}$, yielding the integral representation:
@@
Under this formulation, the conditional distribution of $\bm{x}_{t}$ given the initial state $\bm{x}_{0}\sim q(\bm{x}_{0})$ is Gaussian:
$q(\bm{x}_{t}\mid\bm{x}_{0})=\mathcal{N}\left(\bm{x}_{t};\bm{x}_{0}+\int_{0}^{t}\bm{h}_{s},ds,,t\bm{I}\right).$ (12)
@@
Since the $\bm{h}_{t}$ holds $\int_{0}^{1}\bm{h}_{t}+\bm{x}_{0}=0$, a simplified expression of (13) can be obtained as follows.
$\bm{x}_{t}=(1-t)\bm{x}_{0}+\sqrt{t}\bm{\epsilon}.$ (14)
@@
$d\bm{x}_{t}=\left[f_{t}\bm{x}_{t}-\frac{g_{t}^{2}}{\beta_{t}}\bm{\epsilon}_{t}\right]dt+g_{t},d\bar{\bm{w}}_{t},$ (15)
@@
$\bm{x}_{t-\Delta t}=\bm{x}_{t}+\int_{t}^{t-\Delta t}\bm{h}_{s},ds-\frac{\Delta t}{t}\bm{\epsilon}+\sqrt{\frac{\Delta t(t-\Delta t)}{t}}\widetilde{\bm{\epsilon}},$ (16)
@@
Remark 1.
An important practical insight that arises from (16) is that if the denoising step size is chosen as $\Delta t=t$, then a single-step denoising operation can be performed. In this case, the denoised sample $\bm{x}_{0}$ is obtained in one step as follows.
$\bm{x}_{0}=\bm{x}_{t}-\bm{\epsilon}+0\cdot\widetilde{\bm{\epsilon}}.$ (17)
This enables fast inference, where the diffusion model’s computational latency becomes comparable to that of traditional neural network-based denoisers. In scenarios such as semantic communication, this property is crucial for real-time or low-latency transmission requirements.
@@ IV-B Relationship Between Diffusion Variable and Noisy Receiving
where $\phi=\frac{\mathbb{E}[\|y\|^{2}]-\sigma^{2}}{\gamma\sigma^{2}}$. The following equation can be obtained.
$\frac{\mathbb{E}\left[\|\bm{z}\|^{2}\right]}{\mathbb{E}\left[\|\bm{n}\|^{2}\right]}=\frac{\mathbb{E}\left[\|(1-t)\bm{x}_{0}\|^{2}\right]}{\mathbb{E}\left[\|\sqrt{t}\bm{\epsilon}\|^{2}\right]}.$ (19)
@@
Because $\mathbf{z}$ and $\mathbf{n}$ are uncorrelated, their energies add: $\mathbb{E}[\|\bm{y}\|^{2}]=\mathbb{E}\!\left[\lVert\mathbf{z}\rVert^{2}\right]+\sigma^{2}$. Therefore, the following equation can be obtained.
$\mathrm{SNR}_{\mathrm{obs}}=\frac{\mathbb{E}[\|\bm{y}\|^{2}]-\sigma^{2}}{\sigma^{2}}=\frac{\mathbb{E}\!\left[\lVert\mathbf{z}\rVert^{2}\right]}{\sigma^{2}}=\frac{(1-t)^{2}\gamma}{t}.$ (20)
Introduce $\phi\triangleq\mathrm{SNR}_{\mathrm{obs}}/\gamma>0$; the matching condition can be obtained as follows.
@@
In the training procedure, since the purpose of denoise is to solve $\bm{z}$, thus the $\bm{x}_{0}$ is set equal to $\bm{z}$. Therefore, the $t$ can be simplified as follows.
$t=\frac{2\sigma^{2}+1-\sqrt{1+4\sigma^{2}}}{2\sigma^{2}}.$ (29)
@@
Define the scaled observation $\tilde{\mathbf{y}}=\alpha\mathbf{y}$ with scalar $\alpha>0$. The equality
$\mathbb{E}\!\bigl{[}\lVert\tilde{\mathbf{y}}\rVert^{2}\bigr{]}=\mathbb{E}\!\bigl{[}\lVert\mathbf{x}_{t}\rVert^{2}\bigr{]},$
@@
$\;\alpha=\sqrt{\frac{(1-t)^{2}(\mathbb{E}[\|\bm{y}\|^{2}]-\sigma^{2})+t}{\mathbb{E}[\|\bm{y}\|^{2}]}}=\sqrt{\frac{(1-t)^{2}\gamma+t}{\gamma+\sigma^{2}}}.$ (30)
Moreover, $0<\alpha\leq 1$; $\alpha$ decreases strictly with $t$ and with the observable signal-to-noise ratio $S=(\mathbb{E}[\|\bm{y}\|^{2}]-\sigma^{2})/\sigma^{2}$.
@@
Because $\mathbf{z}$ and $\mathbf{n}$ are uncorrelated, $\mathbb{E}[\lVert\mathbf{y}\rVert^{2}]=\gamma+\sigma^{2}=\mathbb{E}[\|\bm{y}\|^{2}]$. Scaling by $\alpha$ multiplies the second-order moment by $\alpha^{2}$, hence
$\mathbb{E}\!\bigl{[}\lVert\tilde{\mathbf{y}}\rVert^{2}\bigr{]}=\alpha^{2}\mathbb{E}[\|\bm{y}\|^{2}].$ (31)
Equating the above to the target power $(1-t)^{2}\gamma+t$ and substituting $\gamma=\mathbb{E}[\|\bm{y}\|^{2}]-\sigma^{2}$ derived from the observable statistics yields the follows.
$\alpha^{2}\mathbb{E}[\|\bm{y}\|^{2}]=(1-t)^{2}(\mathbb{E}[\|\bm{y}\|^{2}]-\sigma^{2})+t,$ (32)
$\alpha^{2}=\frac{(1-t)^{2}(\mathbb{E}[\|\bm{y}\|^{2}]-\sigma^{2})+t}{\mathbb{E}[\|\bm{y}\|^{2}]}.$ (33)
Taking the principal square root preserves positivity and furnishes the closed form for $\alpha$. Positivity follows from $t\in(0,1)$ and $\mathbb{E}[\|\bm{y}\|^{2}]>\sigma^{2}$. To bound $\alpha$ by unity, observe that $(1-t)^{2}(\mathbb{E}[\|\bm{y}\|^{2}]-\sigma^{2})+t\leq\mathbb{E}[\|\bm{y}\|^{2}]$ because $(1-t)^{2}\leq 1-t$ on $(0,1)$ and $t<1$. Strict decrease in $t$ is evident from the negative derivative of the numerator with respect to $t$; strict decrease in $S$ follows from the monotonic growth of $\mathbb{E}[\|\bm{y}\|^{2}]$ with $\gamma$. ∎
A direct consequence of Theorem 1 and Theorem 2 is that the properly rescaled observation $\tilde{\mathbf{y}}=\alpha\mathbf{y}$ attains the *same second-order distribution* as the synthetic sample $\mathbf{x}_{t}$ employed during diffusion training. Because the clean component of the channel output equals the training datum ($\mathbf{z}=\mathbf{x}_{0}$), the only mismatch between the two random vectors is the pair of linear coefficients that multiplies the signal term and the additive Gaussian term, respectively. Selecting the step size $t$ via Theorem 1 ensures that the signal-to-noise ratio satisfies $(1-t)^{2}\gamma/t=\gamma/\sigma^{2}$; the ensuing scaling factor derived in Theorem 3 simplifies to $\alpha^{2}=t/\sigma^{2}$, so that $\alpha\mathbf{y}=(1-t)\mathbf{x}_{0}+\sqrt{t}\,\hat{\bm{\epsilon}}$ with $\hat{\bm{\epsilon}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Therefore the rescaled received vector replicates exactly the affine stochastic structure of $\mathbf{x}_{t}$. Although $\gamma=\mathbb{E}[\lVert\mathbf{x}_{0}\rVert^{2}]$ appears in both theorems, it is known at training time because it can be replaced by the empirical mean energy $\bar{\gamma}$ computed over the entire corpus of clean samples. In summary, the pair $\{t^{\star},\alpha^{\star}\}$ aligns the second-order statistics—and, when $\mathbf{x}_{0}$ is Gaussian, the full distribution—of the real channel observation with those of the diffusion-generated surrogate, thereby allowing the denoiser trained on $\mathbf{x}_{t}$ to operate on actual channel outputs without retraining or additional calibration.
@@ IV-C LDM Based Sematic Communciation Framework
Given the received latent vector $\tilde{\bm{z}}=\bm{z}+\bm{n}$, where $\bm{z}$ denotes the transmitted latent semantic representation and $\bm{n}\sim\mathcal{N}(0,\sigma^{2}\bm{I})$ represents channel noise, the receiver first estimates the instantaneous SNR. This estimate is used to analytically determine a denoising timestep $t\in(0,1]$ according to Theorem 1, which ensures that the signal–noise energy ratio in the diffusion variable matches that of the channel output. Simultaneously, a linear scaling factor $\alpha$ is computed via Theorem 2 to adjust the received vector’s magnitude, aligning its distribution with that of the diffusion model’s training inputs. This transformation yields the input to the LDM as $\bm{x}_{t}=\alpha\bm{y}$ The LDM then performs reverse diffusion starting from $\bm{\alpha\bm{y}}$, using the analytically determined timestep $t$ from Theorem 1. Different from conventional denoisers, which are often retrained or fine-tuned under new noise conditions, the LDM leverages its learned generative prior to adapt across a wide range of channel perturbations without modification. The reverse process can be executed with a variable number of iterations, depending on the SNR: fewer steps under high-quality channels and more when the received signal is severely corrupted. In particular, the framework supports single-step reverse inference using the approximated expression in (16), which enables low-latency operation when appropriate. Once denoised, the output $\hat{\bm{z}}$ is forwarded to the decoder, which reconstructs the semantic data. Since the LDM operates in latent space, its computational burden is low compared to pixel- or token-level generative models. Furthermore, the use of pretrained LDMs decouples the denoising mechanism from the encoder–decoder training, enabling modular updates. That is, more powerful LDMs—trained on large-scale data or advanced architectures—can be integrated directly, enhancing performance without altering the underlying semantic transceiver structure. This modular, SNR-aware framework allows the semantic communication system to maintain robustness against varying channel conditions while retaining the scalability and efficiency of a latent-space representation. It overcomes the limitations of conventional discriminative models, which suffer from generalization gaps when encountering unseen noise levels or out-of-distribution inputs, by grounding the denoising process in the theoretical structure of stochastic differential equations and distribution alignment.

### v3

@@ III System Model and Problem Formulation
The end-to-end system comprises a source, a joint source–channel encoder, a fading channel with additive white Gaussian noise (AWGN), a denoising module, and a decoder. Let $\bm{x}\in\mathbb{R}^{n}$ denote the original source data. The source encoder, parameterized by $\bm{\phi}$, maps $\bm{x}$ into a compact semantic latent vector $\bm{z}\in\mathbb{R}^{d}$, with $d\ll n$, via
@@
where $f_{\bm{\phi}}:\mathbb{R}^{n}\to\mathbb{R}^{d}$ is typically realized by a variational autoencoder (VAE). The semantic latent vector $\bm{z}$ is transmitted over a fading channel with channel, yielding the received representation
$\tilde{\bm{z}}=\eta\bm{z}+\bm{n},\quad\bm{n}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I}),$ (8)
where $\eta$ is the signal power attenuation factor caused by channel fading, and the $\sigma^{2}$ denotes the channel noise power [^1] Although the analysis in this paper assumes linear channel fading with additive Gaussian noise, [34] demonstrates that a VAE–ADMM pre-processing architecture can transform data corrupted by any non-Gaussian noise into an equivalent representation consisting of amplitude fading followed by zero-mean Gaussian noise. Consequently, the channel model adopted here remains applicable to wireless links whose noise statistics deviate from Gaussian assumptions..
At the receiver, the corrupted latent vector $\tilde{\bm{z}}$ undergoes denoising to mitigate the impact of channel noise. Let $g_{\bm{\psi}}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ represent the denoising function parameterized by $\bm{\psi}$, and $h_{\bm{\theta}}:\mathbb{R}^{d}\to\mathbb{R}^{n}$ denote the semantic decoder parameterized by $\bm{\theta}$. The final reconstructed data $\hat{\bm{x}}$ is obtained as
@@
$\min_{\bm{\phi},\bm{\psi},\bm{\theta}},\mathbb{E}_{\bm{x},\bm{n}}\left[\left\|\bm{x}-h_{\bm{\theta}}\left(g_{\bm{\psi}}(\bm{z}+\bm{n})\right)\right\|_{2}^{2}\right],$ (10)
subject to the encoder constraint $\bm{z}=f_{\bm{\phi}}(\bm{x})$. This formulation captures the full stochasticity of the channel and reflects the end-to-end performance of the semantic transmission process. Two primary challenges arise in solving this problem. First, due to the dimensionality reduction $d\ll n$, the decoder must reconstruct a high-dimensional signal from a compressed and noise-contaminated latent representation. This makes the inverse mapping severely ill-posed and sensitive to noise perturbations. Second, the Gaussian noise introduced during transmission alters the distribution of the latent variable at the receiver, causing a significant mismatch between the training and inference distributions, commonly referred to as the OOD problem. These challenges underscore the need for a robust, distribution-aware denoising mechanism that can generalize across varying SNRs and input distributions. To address this, we propose a theoretically grounded denoising approach based on stochastic differential equations and generative diffusion models, which will be detailed in the next section.
@@ IV-A Theoretical Basis of Diffusion Model-Based Denoising
$d\bm{x}_{t}=f_{t}\bm{x}_{t}dt+g_{t}d\bm{w}_{t},$ (11)
where $\bm{x}_{t}\in\mathbb{R}^{d}$ denotes the latent variable at time $t\in[0,T]$, $\bm{w}_{t}$ is a standard Wiener process, $f_{t}$ is a time-dependent drift coefficient, and $g_{t}$ is the diffusion coefficient. For tractability, we set $g_{t}=1$ and define the drift as $\bm{h}_{t}=f_{t}\bm{x}_{t}$, yielding the integral representation:
@@
[…]
@@
$d\bm{x}_{t}=\left[f_{t}\bm{x}_{t}-\frac{g_{t}^{2}}{\beta_{t}}\bm{\epsilon}_{t}\right]dt+g_{t}d\bar{\bm{w}}_{t},$ (15)
@@
[…]
@@
Remarkably, since the $\bm{h}_{t}$ holds $\int_{0}^{1}\bm{h}_{t}+\bm{x}_{0}=0$, similar to [35] by setting the $\bm{h}_{t}=\bm{x}_{0}$, a simplified expression of (14) and (16) can be obtained as follows .
$\bm{x}_{t}=(1-t)\bm{x}_{0}+\sqrt{t}\bm{\epsilon},$ (17)
$\bm{x}_{t-\Delta t}=(1-\Delta t)\bm{x}_{t}-\frac{\Delta t}{t}\bm{\epsilon}+\sqrt{\frac{\Delta t(t-\Delta t)}{t}}\widetilde{\bm{\epsilon}}.$ (18)
@@ IV-B Relationship Between Diffusion Variable and Noisy Receiving
where $\phi=\frac{\mathbb{E}[\|y\|_{2}^{2}]-\sigma^{2}}{\gamma\sigma^{2}}$. The following equation can be obtained.
$\frac{\mathbb{E}\left[\|\eta\bm{z}\|_{2}^{2}\right]}{\mathbb{E}\left[\|\bm{n}\|_{2}^{2}\right]}=\frac{\mathbb{E}\left[\|(1-t)\bm{x}_{0}\|^{2}\right]}{\mathbb{E}\left[\|\sqrt{t}\bm{\epsilon}\|_{2}^{2}\right]}.$ (20)
@@
Because $\bm{z}$ and $\bm{n}$ are uncorrelated, thus $\mathbb{E}[\|\bm{y}\|_{2}^{2}]=\mathbb{E}[\|\eta\bm{z}+\bm{n}\|_{2}^{2}]=\mathbb{E}\!\left[\lVert\eta\bm{z}\rVert^{2}\right]+\sigma^{2}$. Therefore, the following equation can be obtained.
$\mathrm{SNR}_{\mathrm{obs}}=\frac{\mathbb{E}\!\left[\|\eta\bm{z}\|_{2}^{2}\right]}{\sigma^{2}}=\frac{\mathbb{E}[\|\bm{y}\|_{2}^{2}]-\sigma^{2}}{\sigma^{2}}$ (21)
The purpose of transforming (21) is because, for the receiver, it cannot know the $\|\eta\bm{z}\|_{2}^{2}$ for a specific $\bm{z}$, since it doesn’t the $\bm{z}$, but can only obtain the power of $\|\bm{y}\|_{2}^{2}$ and $\|\bm{n}\|_{2}^{2}$ through measurement. Therefore, it needs to calculate SNR through the above transformation. By introducing $\phi\triangleq\mathrm{SNR}_{\mathrm{obs}}/\gamma>0$, the matching condition of (20) can be obtained as follows.
@@
In the training procedure, since the purpose of denoise is to solve $\bm{z}$, thus if the channel fading factor $\eta$ can be obtained through some channel estimation methods, the $\bm{x}_{0}$ is set equal to $\bm{z}$. Therefore, the $t$ can be simplified as follows.
$t=\frac{2\sigma^{2}+\eta-\sqrt{\eta^{2}+4\eta^{2}\sigma^{2}}}{2\sigma^{2}}.$ (30)
@@
Define the scaled observation $\tilde{\bm{y}}=\alpha\bm{y}$ with scalar $\alpha>0$. The equality
$\mathbb{E}\!\bigl{[}\lVert\tilde{\bm{y}}\rVert^{2}\bigr{]}=\mathbb{E}\!\bigl{[}\lVert\bm{x}_{t}\rVert^{2}\bigr{]},$
@@
$\;\alpha=\sqrt{\frac{(1-t)^{2}(\mathbb{E}[\|\bm{x}_{0}\|_{2}^{2}])+t}{\mathbb{E}[\|\bm{y}\|_{2}^{2}]}}.$ (31)
Moreover, $0<\alpha\leq 1$; $\alpha$ decreases strictly with $t$ [^2] Since $\bm{x}_{0}$ corresponds to the label in the DM training process, its distribution can be obtained..
@@
The expected value of $\mathbb{E}\left[\|\alpha\bm{y}\|^{2}\right]$ is as follows.
$\mathbb{E}\left[\|\alpha\bm{y}\|^{2}\right]$ $=\mathbb{E}\left[\|\alpha(\eta\bm{z}+\bm{n})\|^{2}\right],$
$=\alpha^{2}\left(\mathbb{E}\left[\|\bm{s}\|^{2}\right]+\mathbb{E}\left[\|\bm{n}\|^{2}\right]\right).$
Moreover, the expected value of $\mathbb{E}\left[\|\bm{x}_{t}\|^{2}\right]$ is as follows.
$\mathbb{E}\left[\|\bm{x}_{t}\|^{2}\right]$ $=\mathbb{E}\left[\|(1-t)\bm{x}_{0}+\sqrt{t}\bm{\epsilon}\|^{2}\right],$
$=(1-t)^{2}\mathbb{E}\left[\|\bm{x}_{0}\|\right]+t\mathbb{E}\left[\|\bm{\epsilon}\|^{2}\right],$
$=(1-t)^{2}\mathbb{E}\left[\|\bm{x}_{0}\|\right]+t$
Then, by solving the following equation, the value of $\alpha$ can be obtained as Theorem 2.
$\mathbb{E}\left[\|\bm{x}_{t}\|^{2}\right]=\mathbb{E}\left[\|\alpha\bm{y}\|^{2}\right].$
Taking the principal square root preserves positivity and furnishes the closed form for $\alpha$. Positivity follows from $t\in(0,1)$ and $\mathbb{E}[\|\bm{y}\|_{2}^{2}]>\sigma^{2}$. To bound $\alpha$ by unity, observe that $(1-t)^{2}(\mathbb{E}[\|\bm{y}\|_{2}^{2}]-\sigma^{2})+t\leq\mathbb{E}[\|\bm{y}\|_{2}^{2}]$, because $(1-t)^{2}\leq 1-t$ on $(0,1)$ and $t<1$. Strict decrease in $t$ is evident from the negative derivative of the numerator with respect to $t$. ∎
A direct consequence of Theorem 1 and Theorem 2 is that the properly rescaled observation $\tilde{\bm{y}}=\alpha\bm{y}$ attains the same second-order distribution as the synthetic sample $\bm{x}_{t}$ employed during diffusion training. Because the clean component of the channel output equals the training data pairs as $\bm{z}=\bm{x}_{0}$, the only mismatch between the two random vectors is the pair of linear coefficients that multiply the signal term and the additive Gaussian term, respectively. Selecting the step size $t$ via Theorem 1 ensures that the SNR satisfies $(1-t)^{2}\gamma/t=\gamma/\sigma^{2}$; the scaling factor derived in Theorem 2 ensure the $\mathbb{E}\left[\alpha\bm{y}\right]=\mathbb{E}\left[(1-t)\bm{x}_{0}+\sqrt{t}\bm{\epsilon}\right]$ with $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$. Therefore, the rescaled received vector replicates exactly the affine stochastic structure of $\bm{x}_{t}$. Although $\gamma=\mathbb{E}[\lVert\bm{x}_{0}\rVert^{2}]$ appears in both theorems, it is known at training time because it can be calculated through the whole training data.
@@ IV-C LDM Based Sematic Communciation Framework
Given the received latent vector $\tilde{\bm{z}}=\eta\bm{z}+\bm{n}$, where $\bm{z}$ denotes the transmitted latent semantic representation and $\bm{n}\sim\mathcal{N}(0,\sigma^{2}\bm{I})$ represents channel noise, the receiver first estimates the instantaneous SNR. This estimate is used to analytically determine a denoising timestep $t\in(0,1]$ according to Theorem 1, which ensures that the signal–noise energy ratio in the diffusion variable matches that of the channel output. Simultaneously, a linear scaling factor $\alpha$ is computed via Theorem 2 to adjust the received vector’s magnitude, aligning its distribution with that of the diffusion model’s training inputs. This transformation yields the input to the LDM as $\bm{x}_{t}=\alpha\bm{y}$ The LDM then performs reverse diffusion starting from $\alpha\bm{\bm{y}}$, using the analytically determined timestep $t$ from Theorem 1. Different from conventional denoisers, which are often retrained or fine-tuned under new noise conditions, the LDM leverages its learned generative prior to adapt across a wide range of channel perturbations without modification. The reverse process can be executed with a variable number of iterations, depending on the SNR: fewer steps under high-quality channels and more when the received signal is severely corrupted. In particular, the framework supports single-step reverse inference using the approximated expression in (16), which enables low-latency operation when appropriate. Once denoised, the output $\hat{\bm{z}}$ is forwarded to the decoder, which reconstructs the semantic data. Since the LDM operates in latent space, its computational burden is low compared to pixel- or token-level generative models. Furthermore, the use of pretrained LDMs decouples the denoising mechanism from the encoder–decoder training, enabling modular updates. That is, more powerful LDMs—trained on large-scale data or advanced architectures—can be integrated directly, enhancing performance without altering the underlying semantic transceiver structure. This modular, SNR-aware framework allows the semantic communication system to maintain robustness against varying channel conditions while retaining the scalability and efficiency of a latent-space representation. It overcomes the limitations of conventional discriminative models, which suffer from generalization gaps when encountering unseen noise levels or out-of-distribution inputs, by grounding the denoising process in the theoretical structure of stochastic differential equations and distribution alignment.